The speed of the boat in still water is 15km/hr.It can go 30 km upstream and return to down stream in 41/2hrs.find the speed of the stream.
Answer by lwsshak3(11628)

The speed of the boat in still water is 15km/hr.It can go 30 km upstream and return to down stream in 41/2hrs.find the speed of the stream.
**
let x=speed of stream
(15+x)=speed of boat downstream
(15-x)=speed of boat upstream
travel time=distance/speed
30/(15+x)+30/(15-x)=9/2
60/(15+x)+60/(15-x)=9
LCD: (15+x)(15-x)
60(15-x)+60(15+x)=9(15+x)(15-x)
60(15-x+15+x)=9(225-x^2)
60(30)=2025-9x2
9x^2=2025-1800
9x2=225
take sqrt of both sides
3x=15
x=5
speed of stream=5 km/hr